OpenStudy (anonymous):

PLEASES HELP and show the steps :( Solve the equation. Check the solution. 15/x+9x-7/x+2=9

OpenStudy (anonymous):

Its just a simple equation if you know each simple steps. 15/X + 9X-7/x+2=9 The best and easy way is to cross multtiply. The equation will become- 15(x+2)+9x-7(x)=(x)(x+2}(9) 15X+30+9x^2-7x=(x^2+2X}(9) 15x+30+9X^2-7x=9x^2+18x collecting like terms- 9x^2-9x^2+15x-7x-18x=-30 15x-25x=-30 -10x=-30 x=-30/-10 x=3

OpenStudy (anonymous):

hope this helps!

OpenStudy (anonymous):

Thankyou so much, Shawnee !

OpenStudy (anonymous):

your welcome & no problem ! :)